# When Context Is Not Enough: Modeling Unexplained Variability in Car-Following Behavior

### v1

@@ 3.2 Motivations and Theoretical Background
    Temporal correlations, where errors at adjacent time steps are statistically dependent (Zhang and Sun,, 2024; Zhang et al., 2024b, ).
@@
With the insights from the above discussion, we adopt the formulation proposed in Zhang and Sun, (2024) and Zhang et al., 2024b , representing the action $a_{t}=a(\bm{x}_{t},t)$ of a following vehicle at time $t$ as
@@
[…]
@@
A comparative overview of how the literature incorporates temporal correlation into car-following models is presented in Table 1. Treiber et al., (2006) introduced an OU process to model $\delta(t)$, which provides a continuous-time, mean-reverting formulation. However, the OU process is inherently linear and governed by a fixed decay rate, which imposes a rigid structure on temporal correlations and limits its ability to capture adaptive or nonlinear behaviors observed in real driving contexts. Hoogendoorn and Hoogendoorn, (2010) eliminate $\delta(t)$ based on a Cochrane-Orcutt correction by assuming a first-order autoregressive (i.e., AR(1)) structure in the residuals. While computationally efficient, this approach also assumes a fixed correlation structure and linear dependence, constraining its capacity to model the complex, scenario-dependent dynamics that characterize real-world car-following behavior. More recent efforts, such as Zhang and Sun, (2024), used GPs with stationary kernels to capture smooth and consistent patterns over time. This approach improves flexibility by enabling smooth, nonparametric modeling of correlations. Nonetheless, the assumption of stationarity still imposes a uniform correlation structure over time, which is inadequate for modeling nonstationary phenomena such as sudden braking or transitions between driving regimes. In a related direction, Zhang et al., 2024b adopted higher-order AR processes to capture both positive and negative temporal dependencies, offering improved expressiveness compared to AR(1) and OU processes, but remaining fundamentally limited by their stationary nature.
@@
To provide a comprehensive understanding of our modeling approach, it is beneficial to distinguish the different forms of uncertainty encountered in traffic flow modeling. Broadly, uncertainty can be categorized into two primary groups (Zhou et al.,, 2025): (i) aleatoric uncertainty, reflecting inherent randomness in driving behaviors such as variability in reaction times and driver aggressiveness; and (ii) epistemic uncertainty, arising from limited knowledge, incomplete data, or model mis-specification, such as unobserved driver intentions or omitted explanatory variables. Within these broad categories, the primary focus of our approach, i.e., residual temporal correlation, is conceptually a structured form of aleatoric uncertainty, capturing the memory effects and sequential dependencies in driver decisions not explained by the mean dynamics alone. However, it also overlap with epistemic uncertainty, as persistent patterns in residuals may indicate missing explanatory variables or latent behavioral mechanisms. Our framework explicitly addresses this structured aleatoric uncertainty through a nonstationary Gaussian process kernel, thus enhancing the realism and interpretability of stochastic simulations. Explicit modeling of epistemic uncertainty remains a valuable direction for future research, potentially through hierarchical or latent variable approaches.
@@
Treiber et al., (2006) IDM Ornstein-Uhlenbeck (OU) processes
Hoogendoorn and Hoogendoorn, (2010) GHR/IDM Cochrane-Orcutt correction (i.e., AR(1) process)
Zhang and Sun, (2024) IDM Gaussian processes
Zhang et al., 2024b IDM AR processes with higher orders
@@ 3.3 Temporal Dependencies Modeling in Car-Following Behaviors
To explicitly model the structured temporal dependencies beyond those captured by the RNN, it is crucial to account for the inherent nonlinear dependencies in the data, both within and across temporal observations. Here we adopt a Gaussian processes-based approach (see Zhang and Sun, (2024) for more details) to capture these dependencies, which provides a richer and more realistic representation than a simple identity covariance matrix from the inappropriate i.i.d. assumption, yielding the hybrid formulation:
@@
where $\delta_{\mathrm{GP}}(t;\bm{\lambda})\sim\mathcal{GP}(0,k(t,t^{\prime};\bm{\lambda}))$ is a zero-mean GP governed by a kernel function $k$, parameterized by $\bm{\lambda}$. This formulation augments the neural model with a temporally correlated residual process, effectively decoupling the modeling of nonlinear mean dynamics from that of structured uncertainty. Over a finite time horizon of $T$ steps with sampling interval $\Delta t$, the GP prior over the sequence of temporal residuals $\bm{\delta}_{\mathrm{GP}}=\left[\delta_{\mathrm{GP}}(\Delta t),\ldots,\delta_{\mathrm{GP}}(T\Delta t)\right]^{\top}$ is given by a zero-mean Gaussian distribution $\bm{\delta}_{\mathrm{GP}}\sim\mathcal{N}(\bm{0},\bm{K})$, where $\bm{K}\in\mathbb{R}^{T\times T}$ is the kernel matrix (also known as Gram matrix) with entries $[\bm{K}]_{ij}=k(i\Delta t,j\Delta t;\bm{\lambda})$.
Inspired by the “better batch” training strategy proposed by Zheng et al., (2024), we consider Equation 3 in a vector form over a mini-batch of $\Delta T$ consecutive time steps. The model’s likelihood over the mini-batch can then be written compactly as:
@@ 3.4.1 Stationary kernel: Squared exponential (SE)
The SE kernel assumes that the latent function is infinitely differentiable, making it suitable for modeling smooth, continuous processes. However, this assumption may be too restrictive for physical systems that exhibit abrupt transitions, such as sudden braking or acceleration events in car-following dynamics. In such cases, the SE kernel can lead to over-smoothed predictions that fail to capture rapid behavioral changes.
@@ 3.4.3 Nonstationary kernel: Gibbs
While stationary kernels assume that temporal correlations are invariant over time, real-world driving behavior often exhibits nonstationary dynamics, drivers’ reactions and uncertainties vary across different scenarios. To model such behavior, we adopt the Gibbs kernel, a nonstationary generalization of the SE kernel introduced by Gibbs, (1997). In this formulation, both the variance $\sigma^{2}(t)$ and the lengthscale $\ell(t)$ are treated as input-dependent functions:
@@
where $\bm{\lambda}(t)=[\ell(t),\sigma(t)]$ are parameterized by a neural network conditioned on contextual inputs. This kernel effectively captures both the context-dependent smoothness and heteroskedasticity in the residual process. Following Zheng et al., (2024), an efficient implementation of the Gibbs kernel over a mini-batch of time steps $t\in\{\Delta t,2\Delta t,\dots,\Delta T\Delta t\}$ can be constructed as
@@ 3.5 Neural Network Parameterization for Probabilistic Modeling
To capture both the mean dynamics and temporal uncertainty structure of car-following behavior, we propose parameterizing the key components of the model, i.e., the mean acceleration, kernel lengthscale, and variance, using neural networks.In particular, we instantiate our approach using the DeepAR framework introduced by Salinas et al., (2020), though the proposed formulation is general and adaptable to other sequential models.
@@
To construct the kernel matrix $\bm{K}$, we proceed as follows: for stationary kernels (e.g., SE or Matérn), we use the batch-averaged parameters ¯$\bar{\ell}_{\mathrm{NN}}=\mathrm{mean}(\bm{\ell}_{\text{NN}}^{\text{batch}})$ and $\bar{\sigma}_{\mathrm{NN}}=\mathrm{mean}(\bm{\sigma}_{\text{NN}}^{\text{batch}})$; for the nonstationary Gibbs kernels, the full sequences $\bm{\ell}_{\text{NN}}^{\text{batch}}$ and $\bm{\sigma}_{\text{NN}}^{\text{batch}}$ are used to construct $\bm{K}$ via the formulation in Equation 18.
@@ 4.1 Dataset and Experimental Setup
To evaluate the performance of the proposed DeepAR-based car-following model, we conduct experiments on the HighD dataset (Krajewski et al.,, 2018), which provides high-resolution vehicle trajectory data captured by drones over German highways. Compared to the NGSIM dataset (Punzo et al.,, 2011) , HighD offers improved reliability due to enhanced data collection and post-processing techniques. The dataset consists of 60 drone-recorded videos, each covering a 420-meter stretch of highway, with detailed annotations for vehicle trajectories, velocities, and accelerations.
To reduce redundancy and enhance computational efficiency, we downsample the original data to $5$ Hz (i.e., $\Delta t=0.2$ s) by uniformly selecting every fifth frame. Following the preprocessing pipeline described in Zhang et al., (2021), we transform the trajectory coordinates to a local vehicle-centric frame.
We focus on vehicle pairs exhibiting sustained car-following behavior, selecting only trajectories with duration $t_{0}>30$ seconds to ensure sufficiently informative behavioral patterns. The resulting dataset is denoted as $\mathcal{D}_{\text{HighD}}$. From this, we extract a test set $\mathcal{D}_{\text{test}}$ comprising 30 randomly chosen pairs with even longer durations ($t_{0}>50$) for qualitative evaluation and visualization. The remaining pairs are split into training and validation sets: $\mathcal{D}_{\text{train}}$ contains $70\%$ of the data, while $\mathcal{D}_{\text{val}}$ holds the remaining $30\%$.
For training the DeepAR model, we employ a one-layer LSTM with 64 hidden units. A linear transformation is applied to the final hidden state to predict distributional parameters. A dropout rate of 0.1 is used to mitigate overfitting. The model is optimized using Adam with a learning rate of $1\times 10^{-3}$, a batch size of 256, and trained for up to 500 epochs. Regularization includes weight decay ($1\times 10^{-3}$) and gradient clipping (maximum norm of 1) to ensure numerical stability. The Sigmoid Linear Unit (SiLU) activation function (Hendrycks and Gimpel,, 2016) is employed in the feedforward layers.
@@
The joint distributions offer further insight into the adaptive behavior learned by the model. For instance, sharp braking behavior (characterized by large negative accelerations) coincides with small gaps and high positive relative speeds, confirming expected physical driving responses. Besides, the learned lengthscale $\ell$ and spacing $s$ reveals a non-monotonic relationship that changes direction around a critical threshold of approximately 20 meters: For small gaps ($s\lesssim 20$ m), a weak negative correlation emerges, where $\ell$ tends to decrease as the gap increases. In very tight headways ($s<10$ m), $\ell$ is slightly elevated—reflecting smoother, more cautious behavior under critical proximity. As spacing grows toward 20 m, drivers become more reactive, resulting in shorter memory and a contraction of $\ell$; For large gaps ($s\gtrsim 20$ m), the trend reverses—a positive correlation appears, with $\ell$ gradually increasing alongside $s$. In these wider spacings, drivers adopt more stable and anticipatory control strategies, prompting the model to expand its temporal correlation horizon. These findings suggest that the learned kernel dynamically adapts its memory scale based on the perceived driving context. By presenting both marginals and joint densities, this figure emphasizes how $\ell$ co‐varies with traffic states and highlights the model’s ability to learn a concise, context‐dependent correlation horizon from the training data.
@@ 4.2.1 Simulation formulation
To emulate realistic car-following behavior under uncertainty, we adopt a stochastic simulation framework based on GP with RNN dynamics. This formulation allows us to capture both the mean trajectory and temporally correlated variability in acceleration decisions, producing simulation rollouts that reflect the inherent randomness of human driving. Further technical background on this formulation is available in Zhang and Sun, (2024); Zhang et al., 2024b .
@@
Figure 2 provides a detailed visualization of the car-following dynamics in a 100-vehicle platoon simulation using Algorithm 1. Figure 2(a) shows the individual trajectories of speed, gap, and relative speed over 1000 seconds. Each curve corresponds to a vehicle, the color indicating its position in the platoon. The leader initiates a smooth trapezoidal acceleration-deceleration maneuver, and the system’s response is observed across the following vehicles. Figure 2(b) presents the corresponding time–space diagram for the first 400 seconds, where the vehicle position is plotted against time and colored by instantaneous speed. Following the leader’s maneuver, a disturbance propagates through the follower vehicles. While oscillations emerge in the speed and gap trajectories, their amplitude does not exhibit significant growth along the platoon. In particular, relative speed fluctuations remain bounded across vehicles, and the time–space diagram shows that the speed wave does not intensify as it travels downstream. These observations suggest that the model exhibits marginal string stability under the given scenario. It is important to note that our primary objective is not to guarantee string stability through controller design. Rather, the focus of this work is to develop a data-driven framework that captures stochastic car-following behaviors with realistic temporal patterns. Nonetheless, the observed results indicate that the proposed model maintains acceptable stability properties in multi-vehicle simulations, further supporting its applicability to large-scale traffic modeling.

### v2

@@ 3.2 Motivations and Theoretical Background
    Temporal correlations, where errors at adjacent time steps are statistically dependent (Zhang and Sun, 2024; Zhang et al., 2024b).
@@
With the insights from the above discussion, we adopt the formulation proposed in Zhang and Sun (2024) and Zhang et al. (2024b), representing the action $a_{t}=a(\bm{x}_{t},t)$ of a following vehicle at time $t$ as
@@
where $\bm{x}_{t}=[s_{t},\Delta v_{t},v_{t}]$ denotes the input covariates, $f_{\mathrm{CF}}(\bm{x}_{t})$ as a function of $\bm{x}_{t}$ represents the mean car-following model, $\delta(t)$ accounts for temporal correlations, and $\epsilon_{t}\hbox{\set@color\hskip 5.11108pt\hskip-3.8889pt\hbox{\set@color$\sim$}\hskip-3.8889pt\hskip-5.11108pt\raisebox{6.66875pt}{\hbox{\set@color\tiny{{i.i.d.}}}}\hskip-5.11108pt\hskip 5.11108pt}\,\mathcal{N}(0,\sigma_{0}^{2})$ is an independent and identically distributed (i.i.d.) noise term with variance $\sigma_{0}^{2}$. Note that $\delta(t)$ is the temporally correlated (colored-noise) component, which controls the correlation timescale and smoothness of residual accelerations. The additional term $\epsilon$ captures unstructured observation/estimation noise (e.g., differentiation-induced acceleration noise) and other high-frequency effects, improving identifiability between behavioral temporal dependence and measurement artifacts.
@@
A comparative overview of how the literature incorporates temporal correlation into car-following models is presented in Table 1. Treiber et al. (2006) introduced an OU process to model $\delta(t)$, which provides a continuous-time, mean-reverting formulation. However, the OU process is inherently linear and governed by a fixed decay rate, which imposes a rigid structure on temporal correlations and limits its ability to capture adaptive or nonlinear behaviors observed in real driving contexts. Hoogendoorn and Hoogendoorn (2010) eliminate $\delta(t)$ based on a Cochrane-Orcutt correction by assuming a first-order autoregressive (i.e., AR(1)) structure in the residuals. While computationally efficient, this approach also assumes a fixed correlation structure and linear dependence, constraining its capacity to model the complex, scenario-dependent dynamics that characterize real-world car-following behavior. More recent efforts, such as Zhang and Sun (2024), used GPs with stationary kernels to capture smooth and consistent patterns over time. This approach improves flexibility by enabling smooth, nonparametric modeling of correlations. Nonetheless, the assumption of stationarity still imposes a uniform correlation structure over time, which is inadequate for modeling nonstationary phenomena such as sudden braking or transitions between driving regimes. In a related direction, Zhang et al. (2024b) adopted higher-order AR processes to capture both positive and negative temporal dependencies, offering improved expressiveness compared to AR(1) and OU processes, but remaining fundamentally limited by their stationary nature.
@@
To provide a comprehensive understanding of our modeling approach, it is beneficial to distinguish the different forms of uncertainty encountered in traffic flow modeling. Broadly, uncertainty can be categorized into two primary groups (Zhou et al., 2025): (i) aleatoric uncertainty, reflecting inherent randomness in driving behaviors such as variability in reaction times and driver aggressiveness; and (ii) epistemic uncertainty, arising from limited knowledge, incomplete data, or model mis-specification, such as unobserved driver intentions or omitted explanatory variables. Within these broad categories, the primary focus of our approach, i.e., residual temporal correlation, is conceptually a structured form of aleatoric uncertainty, capturing the memory effects and sequential dependencies in driver decisions not explained by the mean dynamics alone. However, it also overlap with epistemic uncertainty, as persistent patterns in residuals may indicate missing explanatory variables or latent behavioral mechanisms. Our framework explicitly addresses this structured aleatoric uncertainty through a nonstationary Gaussian process kernel, thus enhancing the realism and interpretability of stochastic simulations. Explicit modeling of epistemic uncertainty remains a valuable direction for future research, potentially through hierarchical or latent variable approaches.
@@
Treiber et al. (2006) IDM Ornstein-Uhlenbeck (OU) processes
Hoogendoorn and Hoogendoorn (2010) GHR/IDM Cochrane-Orcutt correction (i.e., AR(1) process)
Zhang and Sun (2024) IDM Gaussian processes
Zhang et al. (2024b) IDM AR processes with higher orders
@@ 3.3 Temporal Dependencies Modeling in Car-Following Behaviors
To explicitly model the structured temporal dependencies beyond those captured by the RNN, it is crucial to account for the inherent nonlinear dependencies in the data, both within and across temporal observations. Here we adopt a Gaussian processes-based approach (see Zhang and Sun (2024) for more details) to capture these dependencies, which provides a richer and more realistic representation than a simple identity covariance matrix from the inappropriate i.i.d. assumption, yielding the hybrid formulation:
@@
where $\delta_{\mathrm{GP}}(t;\bm{\lambda})\sim\mathcal{GP}(0,k(t,t^{\prime};\bm{\lambda}))$ is a zero-mean GP governed by a kernel function $k$, parameterized by $\bm{\lambda}$. The additional term $\epsilon$ is reserved for unstructured, high-frequency effects, effectively acting as a “nugget effect” [^1] The nugget effect refers to the variance of the i.i.d. noise term $\epsilon$, capturing measurement errors and micro-scale variability occurring at scales smaller than the sampling frequency. This ensures that the GP kernel prioritizes structured behavioral dependencies rather than high-frequency sensor noise. commonly used in GP modeling. This formulation augments the neural model with a temporally correlated residual process, effectively decoupling the modeling of nonlinear mean dynamics from that of structured uncertainty. Over a finite time horizon of $T$ steps with sampling interval $\Delta t$, the GP prior over the sequence of temporal residuals $\bm{\delta}_{\mathrm{GP}}=\left[\delta_{\mathrm{GP}}(\Delta t),\ldots,\delta_{\mathrm{GP}}(T\Delta t)\right]^{\top}$ is given by a zero-mean Gaussian distribution $\bm{\delta}_{\mathrm{GP}}\sim\mathcal{N}(\bm{0},\bm{K})$, where $\bm{K}\in\mathbb{R}^{T\times T}$ is the kernel matrix (also known as Gram matrix) with entries $\bm{K}]_{ij}=k(i\Delta t,j\Delta t;\bm{\lambda})$.
Inspired by the “better batch” training strategy proposed by Zheng et al. (2024), we consider Equation 3 in a vector form over a mini-batch of $\Delta T$ consecutive time steps. The model’s likelihood over the mini-batch can then be written compactly as:
@@ 3.4.1 Stationary kernel: Squared exponential (SE)
The SE kernel assumes that the latent function is infinitely differentiable, making it suitable for modeling smooth, continuous processes. However, this smoothness assumption can be restrictive in systems exhibiting rapid regime transitions (e.g., short corrective braking/acceleration bursts), where the effective correlation timescale changes markedly even though acceleration remains physically continuous. In such cases, the SE kernel can lead to over-smoothed predictions that fail to capture rapid behavioral changes.
@@ 3.4.3 Nonstationary kernel: Gibbs
While stationary kernels assume that temporal correlations are invariant over time, real-world driving behavior often exhibits nonstationary dynamics, drivers’ reactions and uncertainties vary across different scenarios. To model such behavior, we adopt the Gibbs kernel, a nonstationary generalization of the SE kernel introduced by Gibbs (1997). In this formulation, both the variance $\sigma^{2}(t)$ and the lengthscale $\ell(t)$ are treated as input-dependent functions:
@@
where $\bm{\lambda}(t)=[\ell(t),\sigma(t)]$ are parameterized by a neural network conditioned on contextual inputs. This kernel effectively captures both the context-dependent smoothness and heteroskedasticity in the residual process. Following Zheng et al. (2024), an efficient implementation of the Gibbs kernel over a mini-batch of time steps $t\in\{\Delta t,2\Delta t,\dots,\Delta T\Delta t\}$ can be constructed as
@@ 3.5 Neural Network Parameterization for Probabilistic Modeling
To capture both the mean dynamics and temporal uncertainty structure of car-following behavior, we propose parameterizing the key components of the model, i.e., the mean acceleration, kernel lengthscale, and variance, using neural networks. In this paper we use DeepAR (Salinas et al., 2020) to instantiate the mean model $f_{\mathrm{CF}}(\cdot)$, but the stochastic residual component $\delta(t)$ is model-agnostic and can be combined with alternative car-following formulations.
@@
To construct the kernel matrix $\bm{K}$, we proceed as follows: for stationary kernels (e.g., SE or Matérn), we use the batch-averaged parameters $\bar{\ell}_{\mathrm{NN}}=\mathrm{mean}(\bm{\ell}_{\text{NN}}^{\text{batch}})$ and $\bar{\sigma}_{\mathrm{NN}}=\mathrm{mean}(\bm{\sigma}_{\text{NN}}^{\text{batch}})$; for the nonstationary Gibbs kernels, the full sequences $\bm{\ell}_{\text{NN}}^{\text{batch}}$ and $\bm{\sigma}_{\text{NN}}^{\text{batch}}$ are used to construct $\bm{K}$ via the formulation in Equation 18.
@@ 4 Stochastic Simulations and Empirical Evaluations
Figure 1: Traffic shockwave visualization of HighD dataset (Track #25, Lane #4), colored by speed (m/s).
@@ 4.1 Dataset and Experimental Setup
To evaluate the performance of the proposed DeepAR-based car-following model, we conduct experiments on the HighD dataset (Krajewski et al., 2018), which provides high-resolution vehicle trajectory data captured by drones over German highways. Compared to the NGSIM dataset (Punzo et al., 2011), HighD offers improved reliability due to enhanced data collection and post-processing techniques. The dataset consists of 60 drone-recorded videos, each covering a 420-meter stretch of highway, with detailed annotations for vehicle trajectories, velocities, and accelerations. Here we choose HighD because acceleration-based stochastic modeling is sensitive to measurement noise, and HighD provides high-quality trajectories with widely adopted preprocessing conventions. While HighD is a freeway dataset, it contains multiple traffic states, including close-following and stop-and-go episodes (as shown in Figure 1), which are simple but sufficient to evaluate regime-dependent temporal correlation.
To reduce redundancy and enhance computational efficiency, we downsample the original data to $5$ Hz (i.e., $\Delta t=0.2$ s) by uniformly selecting every fifth frame. Following the preprocessing pipeline described in Zhang et al. (2021), we transform the trajectory coordinates to a local vehicle-centric frame.
To emphasize challenging interactions within HighD, we include sustained close-following episodes ($t_{0}>30$ s) and evaluate performance across a range of traffic states, including stop-and-go behavior. The resulting dataset is denoted as $\mathcal{D}_{\text{HighD}}$. From this, we extract a test set $\mathcal{D}_{\text{test}}$ comprising 30 randomly chosen pairs with even longer durations ($t_{0}>50$) for qualitative evaluation and visualization. The remaining pairs are split into training and validation sets: $\mathcal{D}_{\text{train}}$ contains $70\%$ of the data, while $\mathcal{D}_{\text{val}}$ holds the remaining $30\%$.
For training the DeepAR model, we employ a one-layer LSTM with 64 hidden units. A linear transformation is applied to the final hidden state to predict distributional parameters. A dropout rate of 0.1 is used to mitigate overfitting. The model is optimized using Adam with a learning rate of $1\times 10^{-3}$, a batch size of 256, and trained for up to 500 epochs. Regularization includes weight decay ($1\times 10^{-3}$) and gradient clipping (maximum norm of 1) to ensure numerical stability. The Sigmoid Linear Unit (SiLU) activation function (Hendrycks and Gimpel, 2016) is employed in the feedforward layers.
@@
The joint distributions offer further insight into the adaptive behavior learned by the model. For instance, sharp braking behavior (characterized by large negative accelerations) coincides with small gaps and high positive relative speeds, confirming expected physical driving responses. Besides, the learned lengthscale $\ell$ and spacing $s$ reveal a non-monotonic relationship that changes direction around a critical threshold of approximately 20 meters. For small gaps ($s\lesssim 20$ m), a weak negative correlation emerges, where $\ell$ tends to decrease as the gap increases. In very tight headways ($s<10$ m), $\ell$ is slightly elevated—reflecting smoother, more cautious behavior under critical proximity. As spacing grows toward 20 m, drivers become more reactive, resulting in shorter memory and a contraction of $\ell$; for large gaps ($s\gtrsim 20$ m), the trend reverses—a positive correlation appears, with $\ell$ gradually increasing alongside $s$. In these wider spacings, drivers adopt more stable and anticipatory control strategies, prompting the model to expand its temporal correlation horizon. These findings suggest that the learned kernel dynamically adapts its memory scale based on the perceived driving context. By presenting both marginals and joint densities, this figure emphasizes how $\ell$ co‐varies with traffic states and highlights the model’s ability to learn a concise, context‐dependent correlation horizon from the training data.
@@ 4.2.1 Simulation formulation
To emulate realistic car-following behavior under uncertainty, we adopt a stochastic simulation framework based on GP with RNN dynamics. This formulation allows us to capture both the mean trajectory and temporally correlated variability in acceleration decisions, producing simulation rollouts that reflect the inherent randomness of human driving. Further technical background on this formulation is available in Zhang and Sun (2024); Zhang et al. (2024b).
@@
Figure 3 provides a detailed visualization of the car-following dynamics in a 100-vehicle platoon simulation using Algorithm 1. Figure 3(a) shows the individual trajectories of speed, gap, and relative speed over 1000 seconds. Each curve corresponds to a vehicle, the color indicating its position in the platoon. The leader initiates a smooth trapezoidal acceleration-deceleration maneuver, and the system’s response is observed across the following vehicles. Figure 3(b) presents the corresponding time–space diagram, where the vehicle position is plotted against time and colored by instantaneous speed. Following the leader’s maneuver, a disturbance propagates through the follower vehicles. While oscillations emerge in the speed and gap trajectories, their amplitude does not exhibit significant growth along the platoon. In particular, relative speed fluctuations remain bounded across vehicles, and the time–space diagram shows that the speed wave does not intensify as it travels downstream. These observations suggest that the model exhibits marginal string stability under the given scenario. It is important to note that our primary objective is not to guarantee string stability through controller design. Rather, the focus of this work is to develop a data-driven framework that captures stochastic car-following behaviors with realistic temporal patterns. Nonetheless, the observed results indicate that the proposed model maintains acceptable stability properties in multi-vehicle simulations, further supporting its applicability to large-scale traffic modeling.